how do i solve sqrt(2x+5) = 3
Answer by ewatrrr(24785)

 
Hi,
sqrt(2x+5) = 3 |Squaring both sides of the EQ
2x+5 = 9
x = 2
CHECKING our Answer***
sqrt(2*2+5) = sqrt(9)= 3